OpenStudy (anonymous):

A=P(1+(r/n))^nt isolate r

OpenStudy (mathstudent55):

\(a = P\left( 1 + \dfrac{r}{n} \right) ^{nt} \)

OpenStudy (anonymous):

thats correct. i can plug in values if that will help.

OpenStudy (mathstudent55):

No. First, divide both sides by P.

OpenStudy (anonymous):

ok done

OpenStudy (anonymous):

then what do i do with the exponents?

OpenStudy (mathstudent55):

\(\dfrac{a}{P} = \left( 1 + \dfrac{r}{n} \right) ^{nt}\) Since the right side is raised tot eh nt power, ne we take the nt root of both sides: \( \large \sqrt[nt]{\dfrac{a}{P}} = \sqrt[nt]{\left( 1 + \dfrac{r}{n} \right) ^{nt}}\) \( \large \sqrt[nt]{\dfrac{a}{P}} = 1 + \dfrac{r}{n}\) Now we subtract 1 from both sides: \( \large \sqrt[nt]{\dfrac{a}{P}} -1 = \dfrac{r}{n}\) Now multiply both sides by n: \( \large \large( \sqrt[nt]{\dfrac{a}{P}} -1 \large)n = r\)

OpenStudy (anonymous):

in my class we are studying logarithms and exponential functions. if i isolate r will it give me the same answer as if i solve using logarithms?

OpenStudy (mathstudent55):

yes

OpenStudy (anonymous):

k cool

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

how would i do it using logarithms?